### v3

# Quasielectrostatic eigenmode analysis of anisotropic particles Asaf Farhi1 and Haim Suchowski1,21

Abstract
Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials. Here we analytically derive the quasielectrostatic eigenmodes and eigenpermittivity relations of anistropic spheres and ellipsoids, and present the results for spheres. We show that anisotropic particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy. Our results pave the way for a new generation of directional and multispectral detectors and biomarkers and apply to other fields of physics.
Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability
to provide precise control over the spatial response
and their unique physical properties. A key distinction of subwavelength particles is their strong response, which can occur without gain [1, 2, 3, 4]. Furthermore, by adjusting their geometry, it is often possible to tailor their response at will. In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [5, 6, 7, 8, 9, 10, 11]. While isotropic nanostructures have widespread use [12], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [13, 14].
When particles are composed of anisotropic media, they typically display resonances in unexplored spectral regions and exhibit exotic properties that arise from the directional and multispectral nature of the bulk [15], holding promise for a new generation of anisotropic resonators and biomarkers.
In addition to this class of resonators, a variety of naturally occurring and existing systems are in fact uniaxial particles, ranging from liquid crystal droplets to ferromagnets and ice grains [16, 17, 18, 19, 20, 21]. Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules. [22, 23, 24, 25].
When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime,
Laplace’s equation plays a dominant role in a variety of fields including quasi-electrostatics, gravity, thermal physics, fluid mechanics, and magnetostatics [26, 27, 28, 29, 30, 31, 32]. Analytic solutions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Recently, the emergence of anisotropic materials in optics [13, 44, 45, 14, 46] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [47, 4, 48, 49, 50, 51, 51, 52, 53, 54].
More specifically, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium
oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3) [55, 14, 56, 20]. Very recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [56].
However, while the dipole response of anisotropic particles is well understood [42], the derivation of higher-order modes and eigenpermittivity relations remains unexplored due to the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
The high-order response of such particles is important in two common scenarios: when the source is in proximity to the particle [14, 57] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [58, 59]. Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [14, 60, 61, 62, 56].
Figure 1: Angular distributions of the quasi-electrostatic potential eigenstates for isotropic, uniaxial, and biaxial spheres. The modes of the isotropic sphere are directed along the field/dipole excitation direction [1, 2] whereas the modes of the anisotropic spheres are aligned with the crystal axes and multispectral (colors correspond to frequencies). The second-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to three superimposed modes.
Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivities relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source [1, 2], to allow several degrees of freedom of the permittivities in the different axes. Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds. In addition, it can readily include the temporal dependence of the response by incorporating the time-varying function of the excitation source. Note that whereas previous (semianalytical) works on particles with cartesian anisotropy focused on dielectric materials with nonresonant excitations, we consider the resonant modes with at least one [52, 51]. Our analysis shows that anisotropic particles exhibit novel radiation patterns. While we consider the localized modes, in practice the particles have size and radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns [63]. In addition, we identify axial-eigenpermittivity sum
rules, resulting in resonance splitting in anisotropic particles. We also show that the physics of anisotropic spheres is different from anisotropic ellipsoids: anisotropic spheres support uniaxial modes and exhibit eigenfrequency degeneracy, which leads to coherently superimposed modes, that anisotropic ellipsoids do not support. We apply our theory to spherical uniaxial and biaxial particles composed of hBN and MoO3 phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 64, 65, 66].
Our starting point in calculating the eigenstates in the quasistatic regime is
solving Laplace’s equation without a source, that reads for
anisotropic media [67]
where is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor
inside the inclusion and
in the host medium ( is the identity matrix). Note that the physical permittivity tensor generally depends on and the resonances occur when it approximately satisfies the eigenpermittivity relations. Clearly, have to satisfy
the boundary conditions at the particle interface. Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system. In addition to the challenge that the symmetry of
the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely
As a result, the widely-employed spherical harmonics are in general not eigenstates
of anisotropic spherical particles.
In Fig. 1, left column, we present isotropic, uniaxial, and biaxial particles, which are the most common types of particles.
We now analyze the eigenmodes and eigenpermittivities of anisotropic
spheres and ellipsoids. We start with the dipole response
and proceed to the high-order mode responses.
The anisotropic
dipole eigenstates and eigenpermittivities of a sphere and ellipsoid
can be deduced from the scattering analyses in the literature [42, 36]
and read for a dipole mode oriented along
where in the case of an ellipsoid, are its intersections
with the positive axes, respectively, and the second expression for the ellipsoid is on the external envelope. While these eigenstates
and eigenpermittivities appear to be identical to the isotropic ones,
here the modes are oriented along the resonant crystal axis independently
of the incoming field polarization, even though the polarization direction
does affect the excitation magnitude.
We now proceed to the high-order modes of anisotropic particles. In our analysis we choose to express the
eigenstates in cartesian coordinates due to the crystal symmetry with functions
that are suitable for anisotropic medium i.e., ones that enable when substituted in Laplace’s equation and satisfy boundary conditions.
Since the field outside the sphere can be expanded with spherical harmonics,
the high-order spherical harmonics are valid for isotropic
medium [2, 35],
and there
is continuity of the potential on the particle envelope, we suggest
that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [1, 2].
To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and it is sufficient to require . Interestingly, such functions can be composed of the isotropic-sphere eigenstates Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
It can be seen that since in the boundary conditions and modes composed of and satisfy the isotropic eigenpermittivities [1, 2] for the equal permittivity axes and new eigenpermittivity relations, respectively. However, our modes in both cases are directional as opposed to the isotropic sphere modes.
Note that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere. (a) A uniaxial sphere and the permittivity of bulk hBN. (b) The scattering spectrum and of the dipole modes on an hBN sphere surface. Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively. The crystal axes [100],[010],[001] are directed alfong the axes, respectively. (c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material. (d) Scattering spectrum and of the third-order mode, where is excited at two s. In addition, there are geometric-anisotropic shifts compared to the bulk for all modes. Interestingly, high-order resonances have lower frequencies compared to the dipole modes, unlike isotropic spheres.
In some cases superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of
Thus, we suggest the second-order biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
The
expression above is an interesting and novel eigenpermittivity sum rule,
this time for a general biaxial medium.
In addition, one can see that there is a degeneracy
for uniaxial particles with since both are excited. Note that one could select the basis for the uniaxial modes. In Fig. 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles. It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions, and their direction is determined only by the crystal orientation. Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
Figure 3: Spectrum and eigenmodes of a subwavelength biaxial particle. (a) A biaxial sphere
with the bulk permittivities of . The scattering spectra and of the eigenmodes on a sphere surface for the dipole (b), second-order (c), and third-order (d) modes. Note that there are geometric-anisotropic frequency shifts for all modes and resonance splitting for the second-order modes compared to the bulk.
We now proceed to the third-order biaxial mode of an anisotropic sphere. We continue along the lines of the second-order biaxial mode and write Using the procedure above it can be shown that is the resonance condition. In Fig. 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances. We then show for a uniaxial sphere composed of hBN the spectrum and of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes. Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the second-order modes, which also display a degeneracy.
In Fig. 3 we present a biaxial sphere and the axial permittivities of bulk with three resonances. We then show the scattering spectra and on the sphere surface for the dipole (b), second-order (c), and third-order modes (d). Similarly to the uniaxial particle, the second-order modes exhibit resonance splitting and all the modes display geometric-anisotropic frequency shifts. However, here the degeneracy is lifted since there are three different permittivities and the uniaxial modes are not present. Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, unlike the situation in isotropic spheres.
We now consider the second-order modes for anisotropic ellipsoids in analogy to the anisotropic sphere.
We utilize inside the anisotropic
ellipsoid, similarly to the isotropic case of an ellipsoid in Ref. [36].
To impose continuity of we project in the
direction perpendicular to the interface with the unit vector
Comparing inside the ellipsoid to the isotropic case
we obtain that since and (inside and outside) are the same in both
cases, we have the general sum rule
and for the sphere limit we get as expected. The modes in the form similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive and not likely to be realized experimentally.
We proceed to the third-order mode of an anisotropic ellipsoid. We write the potential and field inside the inclusion
Similarly, we compare between inside the ellipsoid in the isotropic and anisotropic cases and obtain that a resonance occurs when
We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement
We verified our theory for two biaxial ellipsoids composed of with hyperspectral midIR near-field imaging in Ref. [56] and obtained excellent agreement for both the modes and the spectra. In that work the near-field imaging tip scattered the field component in the axial direction to the far field, which was then measured.
In conclusion, we presented a framework to derive the
modes and eigenpermittivity relations of anisotropic spherical and
ellipsoid particles in the quasistatic regime. While isotropic structures
have discrete eigenpermittivities, we showed that anisotropic particles exhibit eigenpermittivity sum rules that couple
the axial permittivities. We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting of anisotropic particles. Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support.
The multispectral behavior of anisotropic particles may allow them to function as a biomarker with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [60].
Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [61, 68, 69]. Furthermore, our derived modes of anisotropic particles are expected to have implications for the radiation patterns in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Our results
apply also to other fields of physics such as heat conduction and quasimagnetostatics [57].
Acknowledgement
E. E. Narimanov, J. Jeffet, and Y. Reches are acknowledged for the insightful comments.
We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.