### v4

# Quasielectrostatic eigenmode analysis of anisotropic particles Asaf Farhi1 and Haim Suchowski1,21

Abstract
Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials. Here we derive closed-form expressions for the quasielectrostatic eigenmodes and eigenpermittivity relations of uniaxial and biaxial spheres and ellipsoids. We show that such particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy. Finally, we analytically predict the Q-factors of the resonance peaks. Our results pave the way for a new generation of directional and multispectral detectors and biomarkers and apply to other fields of physics.
Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability
to provide precise control over the spatial response
and their unique physical properties. A key distinction of subwavelength particles is their strong response, which can occur without gain [1, 2, 3, 4]. Furthermore, by adjusting their geometry, it is often possible to tailor their response at will. In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [5, 6, 7, 8, 9, 10, 11]. While isotropic nanostructures have widespread use [12], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [13, 14].
When particles are composed of anisotropic media, they typically display resonances in unexplored spectral regions and exhibit exotic properties that arise from the directional and multispectral nature of the bulk [15], holding promise for a new generation of anisotropic resonators and biomarkers.
In addition to this class of resonators, a variety of naturally occurring and existing systems are in fact uniaxial particles, ranging from liquid crystal droplets to ferromagnets and ice grains [16, 17, 18, 19, 20, 21]. Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules. [22, 23, 24, 25].
When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime,
Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [26, 27, 28, 29, 30, 31, 32]. Closed-form resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Recently, the emergence of anisotropic materials in optics [13, 44, 45, 14, 46] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [47, 4, 48, 49, 50, 51, 51, 52, 53, 54].
More specifically, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium
oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3,) and various geometries have been investigated [55, 14, 56, 20]. Very recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [56].
However, while previous approaches have successfully addressed particles composed of isotropic materials by deriving closed-form expressions of the eigenstates and resonance conditions, fundamental challenges are encountered when analyzing anisotropic systems, particularly in predicting higher-order mode behavior. These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems. This motivates our exploration of a new theoretical framework.
The high-order response of anisotropic particles is important in two common scenarios: when the source is in proximity to the particle [14, 57] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [58, 59]. Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [14, 60, 61, 62, 56].
Figure 1: Angular distributions of the quasi-electrostatic potential eigenstates for isotropic, uniaxial, and biaxial spheres. The modes of the isotropic sphere are directed along the field/dipole excitation direction [1, 2] whereas the modes of the anisotropic spheres are aligned with the crystal axes and multispectral (colors correspond to frequencies). The high-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to superimposed modes.
Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivities relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source [1, 2], to allow several degrees of freedom of the permittivities in the different axes. Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds. In addition, it can readily include the temporal dependence of the response by incorporating the time-varying function of the excitation source. Note that whereas previous (semianalytical) works on particles with cartesian anisotropy primarly focused on dielectric materials with all axial permittivities positive [63, 51], here we explore the resonant modes that emerge when at least one of the principal components is negative . Our analysis shows that anisotropic particles exhibit novel radiation patterns. While we consider the localized modes, in practice subwavelength particles radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns, see Supplementary Material (SM) for details. Even though the radiation of subwavelength particles in free space via high-order modes is negligible, when there is one or more particles or excitation source at a distance from the particle, these modes play a more significant role in the radiation. In addition, we identify axial-eigenpermittivity sum
rules, resulting in resonance splitting in anisotropic particles. We also show that the physics of anisotropic spheres is different from anisotropic ellipsoids: anisotropic spheres support uniaxial modes and exhibit eigenfrequency degeneracy, which leads to coherently superimposed modes, that anisotropic ellipsoids do not support. We apply our theory to spherical uniaxial and biaxial particles composed of hBN and MoO3 phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 64, 65, 66].
Our starting point in calculating the eigenstates in the quasistatic regime is
solving Laplace’s equation without a source, that reads for
anisotropic media [67]
where is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor
inside the inclusion and
in the host medium ( is the identity matrix). Note that the physical permittivity tensor generally depends on and the resonances occur when it approximately satisfies the eigenpermittivity relations. Clearly, have to satisfy
the boundary conditions at the particle interface. Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition. In addition to the challenge that the symmetry of
the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely
As a result, the widely-employed spherical harmonics are in general not eigenstates
of anisotropic spherical particles.
In Fig. 1, left column, we present isotropic, uniaxial, and biaxial particles, which are the most common types of particles.
We now analyze the eigenmodes and eigenpermittivities of anisotropic
spheres and ellipsoids. We start with the dipole response
and proceed to the high-order mode responses.
The anisotropic
dipole eigenstates and eigenpermittivities of a sphere and ellipsoid
can be deduced from the scattering analyses in the literature [42]
and read for a dipole mode oriented along
where in the case of an ellipsoid, are its intersections
with the positive axes, respectively, and the second expression for the ellipsoid is on the external envelope. While these eigenstates
and eigenpermittivities appear to be identical to the isotropic ones [36],
here the modes are oriented along the resonant crystal axis independently
of the incoming field polarization, even though it
does affect the excitation magnitude.
We now proceed to the high-order modes of anisotropic particles. In our analysis we choose to express the
eigenstates in cartesian coordinates due to the crystal symmetry with functions
that are suitable for anisotropic medium i.e., ones that enable when substituted in Laplace’s equation and satisfy boundary conditions.
Since the field outside the sphere can be expanded with spherical harmonics,
the high-order spherical harmonics are valid for isotropic
medium [2, 35],
and there
is continuity of the potential on the particle envelope, we suggest
that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [1, 2].
To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and it is sufficient to require to satisfy the equation. Interestingly, such functions can be composed of the isotropic-sphere eigenstates Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
It can be seen that since in the boundary conditions and modes composed of and satisfy the isotropic eigenpermittivities [1, 2] for the equal permittivity axes and new eigenpermittivity relations, respectively. However, our modes in both cases are directional as opposed to the isotropic sphere modes. Alternatively, one could select the basis .
Note also that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere. (a) A uniaxial sphere and the permittivity of bulk hBN. (b) The scattering spectrum and of the dipole modes on an hBN sphere surface. Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively. The crystal axes [100],[010],[001] are directed along the axes, respectively. (c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material. (d) Scattering spectrum and of the third-order mode, where is excited at two s. Moreover, there are geometric-anisotropic shifts compared to the bulk for all modes. Interestingly, high-order resonances have lower compared to the dipole modes, unlike isotropic spheres.
In some cases superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of , which have to obey only the boundary condition.
Thus, we suggest the following biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
Note that this class of functions preserves the form under Thus, the axial permittivities satisfy novel and interesting sum rules,
this time for a general biaxial medium.
In Fig. 1 we present the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles. It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions, and their direction is determined only by the crystal orientation. Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
Figure 3: Spectrum and eigenmodes of a subwavelength biaxial particle. (a) A biaxial sphere
with the bulk permittivities of . The scattering spectra and of the eigenmodes on a sphere surface for the dipole (b), second-order (c), and third-order (d) modes. Note that there are geometric-anisotropic frequency shifts for all modes and resonance splitting for the second-order modes compared to the bulk.
Furthermore, for a biaxial material whose bulk exhibits three resonances, our resonance conditions predict 18 different peaks in the spectrum in addition to the 3 dipole resonances. In the SM we show that the derived uniaxial modes are all the modes and if additional biaxial modes exist they are typically negligible.
In Fig. 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances. We then show for a uniaxial sphere composed of hBN the spectrum and of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes. Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the high-order modes, which also display a degeneracy.
In Fig. 3 we present a biaxial sphere and the axial permittivities of bulk with three resonances. We then show the scattering spectra and on the sphere surface for the dipole (b), second-order (c), and third-order modes (d). Similarly to the uniaxial particle, the second-order modes exhibit resonance splitting and all the modes display geometric-anisotropic frequency shifts. However, here the degeneracy is lifted since there are three different permittivities and the uniaxial modes are not present. Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, unlike the situation in isotropic spheres.
We now consider the high-order modes of anisotropic ellipsoids in analogy to the anisotropic sphere. Interestingly, the biaxial-sphere modes inside the inclusion volume correspond to the second type of Niven’s functions used in isotropic ellipsoids, at least up to fourth order [36, 68, 69, 70]
We utilize inside the anisotropic
ellipsoid, similarly to the isotropic case of an ellipsoid in Ref. [36].
To impose continuity of for we project in the
direction perpendicular to the interface with the unit vector
Comparing inside the ellipsoid to the isotropic case
we obtain that since and (inside and outside) are the same, we can equate to get an eigenmode and obtain the general sum rule
and for the sphere limit we get as expected. The modes in the form similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive and not likely to be realized experimentally.
We proceed to the third-order mode of an anisotropic ellipsoid. We write the potential and field inside the inclusion
Similarly, we compare inside the ellipsoid in the isotropic and anisotropic cases and obtain that a resonance occurs when
We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement
In this way one can also find the fourth-order biaxial ellipsoid mode and resonance condition.
Finally, we verified our theory for uniaxial and biaxial spheres and a biaxial ellipsoid in simulations (see Appendix) and also experimentally for two biaxial ellipsoids with hyperspectral midIR near-field imaging in Ref. [56], both with excellent agreement for the modes and the spectra.
In conclusion, we presented a framework to derive the
modes and eigenpermittivity relations of anisotropic spherical and
ellipsoid particles in the quasistatic regime. While isotropic structures
have discrete eigenpermittivities, we showed that anisotropic particles exhibit eigenpermittivity sum rules that couple
the axial permittivities. We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting in anisotropic particles. Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support. Importantly, our derived modes can be used to calculate the eigenstates and resonance conditions of clusters of anisotropic particles [35].
The multispectral behavior of anisotropic particles may allow them to function as a biomarker with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [60].
Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [61, 71, 72]. Furthermore, our derived modes of anisotropic particles are expected to have implications for the radiation patterns in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Finally, our results
apply also to other fields of physics such as heat conduction and quasimagnetostatics [57].
Acknowledgement
E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, and Y. Reches are acknowledged for the insightful comments.
We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
I Appendix
Comparison of the theory to simulations
Figure 4: Scattering spectrum, peak widths, and fields near resonance for a uniaxial hBN sphere of a radius excited by an oscillating dipole located at with a dipole moment of . Top: over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement. The circles are the peak widths obtained analytically from compared to the ones calculated from the simulations, with very good agreement. (a),(b),(c),(d),(e) for the resonance conditions: calculated at
respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement. Bottom: over the particle envelope calculated in COMSOL for a dipole located at (same dipole distance as the previous case) with a dipole moment of As predicted the modes disappeared from the spectrum.Figure 5: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 sphere of a radius excited by an oscillating dipole located at with a dipole moment of . Top: calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement. (a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement. Figure 6: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 ellipsoid with excited by an oscillating dipole located at with a dipole moment of . Top: Surface integral over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement. (a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig. 5, defined in the ellipsoid volume).
We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes and the permittivity model used in Figs. 2,3 [55]. The sphere is of radius and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of since we assume [67, 73, 2], and later on prove, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in . This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, see SM for details. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflects the presence of higher spatial frequencies near the dipole, which excite higher-order modes. We then calculate the peak widths analytically and from the simulations. To demonstrate the mode coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis. In this case we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required.
In Fig. 4 we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows in Fig. 4) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t accounted for it). Clearly, there is remarkable agreement between them. This showcases our important ability to analytically predict the Q factors given by
We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes. While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot for a dipole located at the same distance as the previous case and oriented along (31nm). As we predicted the modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles.
We then considered a biaxial subwavelength sphere composed of MoO3 with the permittivity model in Ref. [55]. The sphere is excited by a dipole located at with the dipole moment of
We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL. Here, again, the dipole moment is directed along to preserve the angular dependency of the mode in the mode-excitation coefficient. In Fig. 5 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of . Crucially, our simulations confirm the resonance splitting effect predicted analytically, see arrows in Fig. 5. We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement. We also report very good agreement between the peak frequency widths obtained analytically and the simulations (accurate comparison would require a highly-dense frequency sweep and simulations that are very extensive).
Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with composed of MoO3 with the permittivity model in Ref. [55]. The ellipsoid is excited by a dipole located at with the dipole moment of
We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL. In Fig. 6 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were easily calculated by finding the maxima of . We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement. We note that the isotropic ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following:
As can be seen, for an isotropic ellipsoid with these semi axis ratios and and due to the anisotropic material these resonance are split, as can be seen in Fig. 6. This differs from the case of spheres where the splitting occurs for modes of the same order. Also, in this case the anisotropy of the material could in principle lead to resonance merging.
Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellispoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
II Supplementary Material
II.1 Correspondence between the derived modes and the radiation patterns
Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres. We utilize the TM vector spherical harmonics (VSH). This field pattern is proportional to the sphere mode outside the inclusion [74], which gives the quasistatic results in the limit of Even though this mode is usually considered for isotropic spheres, the results can be applied to the modes of anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs. We write a TM VSH:
where is the spherical Hankel function of the first kind.
Using the vector identity for
we get for the first term:
where
For the second term we use the identity
to get
Then, we utilize the spherical-harmonic identity
to obtain
All in all we get:
We see that when the second term, which is longitudinal, dominates with scaling, and for
the first term, which is transverse, dominates (actually its first term that scales as
Thus, one can conclude from the superposition of that participate in the
quasistatic anisotropic modes, that the radiation pattern will be of the form:
Please note that which gives
overall scaling of the radiated field and we get the following radiation pattern:
II.2 Biaxial sphere modes as superpositions of spherical harmonics
We express the second-order biaxial modes as superpositions of spherical harmonics:
We now proceed to the third and fourth-order biaxial modes. We continue along the lines of the second-order biaxial mode and write Using the procedure above, it can be shown that and are the resonance conditions, respectively. Note that the modes and their resonance conditions readily follow.
II.3 Completeness of the solution
We now prove that the derived uniaxial-sphere modes are all the eigenmodes.
Let us assume that there is an additional eigenstate besides the four modes for each .
Since the isotropic-sphere modes span the space inside the sphere, this eigenstate is a superposition of the remaining isotropic-sphere modes
An eigenstate must satisfy anisotropic Laplace’s equation inside the inclusion, which reads:
Since it is a superposition of isotropic-sphere modes we write:
However, all these modes have orders of that are 2 or higher (including from ) and each function is fundamentally different from the others. Hence, such an eigenstate does not exist, and we have a contradiction. Conclusion: the derived uniaxial sphere modes are all the modes. Note that the uniaxial mode basis does not span space as in the standard case.
Similarly, we write for a biaxial particle
If we get which is a uniaxial medium, and we therefore exclude this case.
All the modes are isotropic (from the boundary conditions) and the modes satisfy when Even if such modes exist, their axial permittivities must satisfy both Laplace’s equation (in a non-trivial manner) and the boundary condition, which imposes strong constraints. As a result, their potential contribution is typically weak. Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
II.4 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
II.4.1 Theory
We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the Appendix.
Dipole mode
The resonance condition is and we therefore find the maximum of
which occurs for hBN [55] at The mode field inside the particle is given by
2nd order modes
For the first 2nd-order mode we get from the theory
which for hBN approximately occurs at We add the degenerate modes, which are excited with equal strengths:
second mode
The derived resonance condition reads
For hBN this is approximately satisfied for
Now we calculate the field by adding the degenerate modes:
3rd order modes
First mode
The resonance condition reads:
For hBN this occurs at
. Adding the degenerate modes we get
To simplify things, we look at the plane
Second mode
The resonance condition is satisfied
for hBN for Adding the modes we obtain:
We plot these field magnitude distributions in Fig. 4.
II.5 Eigenmode expansion for an anisotropic inclusion
Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [74, 67].
Here we expand the quasielectrostatic potential for an anisotropic inclusion.
For concreteness, we consider a uniaxial inclusion and the derived eigenstates. Generalizing it to a biaxial inclusion readily follows.
We start by writing Poisson’s equation for a medium with an anisotropic inclusion:
where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity.
Now we move one term to the lhs in order to be able to define an
eigenvalue equation.
where . The Green function in this case will be for a medium with an inclusion
with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
where we defined to adjust it to an eigenvalue equation with an operator operating on and we performed integration by parts.
We express the electric potential and expand it using the anisotropic
eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
Now we expand using the eigenstates of a particle with anisotropy in the axis:
where is the electric potential for the dipole in free space.
All in all we get:
One can express the eigenvalue in the denominator in terms of the physical parameter using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs. 2 and 3). Note that the first sum includes only the dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper. It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.
Note that in Fig. 4 there is a very small peak at This peak may be due to an off-resonance uniaxial mode as could be understood from the above expansion when applied to biaxial inclusions. Such an expansion is composed of three sums where the second one corresponds to the off resonance uniaxial-mode contributions. This qualitatively agrees with the potential calculated from the field in the simulation results of the form similarly to the mode constructed by adding the modes. Interestingly, this very small peak is not present in the spectrum of the biaxial ellipsoid, in agreement with our prediction that ellipsoids do not favor uniaxial modes.