### v5

# Universal framework for anisotropic particles with resonance laws and splitting

Abstract
Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and richness of mode responses in subwavelength structures. Anisotropic subwavelength particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions. Here, we establish a universal quasistatic framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles. Closed-form solutions reveal axial-permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns. Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss. Full-wave simulations of h-BN and -MoO3 nanoparticles confirm the theory and highlight its predictive power. This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, opening pathways to a new generation of photonic devices with tunable multispectral response and controlled emission patterns, including directional detectors and spectrally selective mid infrared sensors.
Main
Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability to provide precise control over the spatial response In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [1, 2, 3, 4, 5, 6, 7]. While isotropic nanostructures have widespread use [8], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [9, 10]. When particles are composed of anisotropic media, they typically display resonances in unexplored spectral regions and exhibit exotic properties that arise from the directional and multispectral nature of the bulk [11], holding promise for a new generation of anisotropic resonators and biomarkers. In addition to this class of resonators, a variety of naturally occurring and existing systems are in fact uniaxial particles, ranging from liquid crystal droplets to ferromagnets and ice grains [12, 13, 14, 15, 16, 17]. Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules [18, 19, 20, 21].
When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime, Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [22, 23, 24, 25, 26, 27, 28]. Analytic resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Remarkably, to date, the only structure in physics for which closed-form quasistatic eigenmodes and eigenvalues have been derived is the subwavelength isotropic sphere [40, 41, 31]. Recently, the emergence of anisotropic materials in optics [9, 42, 43, 10, 44] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [45, 46, 47, 48, 49, 50, 50, 51, 52, 53]. Moreover, uniaxial and biaxial materials have been shown to be very common in semiconductors and crystals, including the following: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3), with various geometries investigated [54, 10, 55, 16]. Very recently, an experiment on the modes and resonances of biaxial nanoparticles was performed for the first time by the authors and their colleagues [55]. However, while previous approaches have successfully addressed particles composed of isotropic materials by deriving closed-form expressions of the eigenstates and resonance conditions, fundamental challenges are encountered when analyzing anisotropic systems, particularly in predicting higher-order mode behavior. These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems. This motivates our exploration of a new theoretical framework. The high-order response of anisotropic particles is important in two common scenarios: when the source is in proximity to the particle [10, 56] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [57, 58]. Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [10, 59, 60, 61, 55].
Here, we present the second-ever closed-form eigenmode and eigenpermittivity solution for a structure, derived for uniaxial and biaxial spheres embedded in an isotropic medium. While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source [41, 31], to allow several degrees of freedom of the permittivities in the different axes. We identify axial-eigenpermittivity sum rules, leading to resonance splitting in anisotropic particles. In addition, our analysis shows that anisotropic spheres exhibit novel radiation patterns. Furthermore, we generalize our results to uniaxial and biaxial ellipsoids, obtaining the corresponding eigenpermittivity sum rules. Moreover, we analytically derive the Q factors for such anisotropic particles. We apply our theory to uniaxial and biaxial particles composed of hBN and MoO3 phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [62, 63, 64, 65, 55]. Finally, we compare our results to full-wave simulations and find excellent agreement. Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds. In addition, it can readily include the temporal dependence of the response by incorporating the time-varying function of the excitation source. Note that whereas previous (semianalytical) works on particles with cartesian anisotropy primarly focused on dielectric materials with all axial permittivities positive [66, 50], here we explore the resonant modes that emerge when at least one of the principal components is negative . While we consider the localized modes, in practice subwavelength particles radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns, as we derive in Methods. Even though the radiation of subwavelength particles excited by a far field in free space via high-order modes is negligible, when there is one or more particles or excitation source at a distance from the particle, these modes play a significant role in the radiation.
Results
Our starting point in calculating the eigenstates of uniaxial and biaxial anisotropic spheres in the quasistatic regime is solving Laplace’s equation without a source, which reads for anisotropic media [67]
where is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor inside the inclusion and in the host medium ( is the identity matrix). Note that the physical permittivity tensor generally depends on and the resonances occur when it approximately satisfies the eigenpermittivity relations. Clearly, have to satisfy the boundary conditions at the particle interface. Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition. In addition to the challenge that the symmetry of the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely As a result, the widely-employed spherical harmonics are in general not eigenstates of anisotropic spherical particles. In Fig. 1, left column, we present isotropic, uniaxial, and biaxial spheres, which are the most common types of spheres.
We now analyze the eigenmodes and eigenpermittivities of such anisotropic spheres. We start with the dipole response and proceed to the high-order mode responses. The anisotropic dipole eigenstates and eigenpermittivities of a sphere can be deduced from the scattering analyses in the literature [37] and read for a dipole mode oriented along
While these eigenstates and eigenpermittivities appear to be identical to the isotropic ones [37], here the modes are oriented along the resonant crystal axis independently of the incoming field polarization, even though the polarization direction does affect the excitation magnitude.
We now proceed to the high-order modes of anisotropic spheres. In our analysis we choose to express the eigenstates in cartesian coordinates due to the crystal symmetry with functions that are suitable for anisotropic medium i.e., ones that enable when substituted in Laplace’s equation and satisfy boundary conditions. Since the field outside the sphere can be expanded with spherical harmonics, the high-order spherical harmonics are valid for isotropic medium [31, 32], and there is continuity of the potential on the particle envelope, we suggest that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [41, 31]. To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and they naturally satisfy the above requirement. Interestingly, such functions can be composed of the isotropic-sphere eigenstates Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
where correspond to and the first condition holds if either or , whereas the second holds only for resulting in resonance splitting for (two closely-spaced resonances instead of one for isotropic spheres). One can see that since in the boundary conditions and modes composed of and satisfy the isotropic eigenpermittivities [41, 31] for the equal permittivity axes and new eigenpermittivity relations, respectively. However, our modes in both cases are directional as opposed to the isotropic sphere modes. Alternatively, one could select the basis for uniaxial spheres. Note also that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
In some cases, superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of , which have to obey only the boundary condition. Thus, we derive the following biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
These eigenstates are superpositions of spherical harmonics in all space, as is explained in Methods. Note that when imposing the boundary condition, this class of functions preserves the functional form under inside the particle. Thus, the axial permittivities satisfy novel and interesting sum rules, this time for a general biaxial medium. These sum rules result in resonance splitting compared to the isotropic-sphere resonances: anisotropic bulk media usually exhibit a distinct resonance in each axis. Thus, for a resonance in a given axis in the bulk medium (), usually the sums result in two resonances occurring at different frequencies. Hence, for a biaxial material whose bulk exhibits three resonances, our resonance conditions predict 18 different peaks in the spectrum in addition to the 3 dipole resonances. In the Supplementary Material (SM) we show that the derived uniaxial modes form a complete set and if additional biaxial modes exist they are typically negligible. In Fig. 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles. One can see that anisotropic particles support significantly different modes, which typically appear in several spectral regions as each axial permittivity experiences a resonanance at a different frequency (colors correspond to frequencies). In addition, their direction is determined only by the crystal orientation. Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
In Fig. 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances. We then show for a uniaxial sphere composed of hBN the spectrum and of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes. Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the high-order modes, which also display a degeneracy. In Fig. 2 (e) we show in the radiation patterns of the uniaxial sphere excited by an oscillating dipole (Methods). In Fig. 3 we present a biaxial sphere and the axial permittivities of bulk with three resonances. We then show the scattering spectra and on the sphere surface for the dipole (b), second-order (c), and third-order modes (d). Similarly to the uniaxial particle, the second-order modes exhibit resonance splitting and all the modes display geometric-anisotropic frequency shifts. However, here the degeneracy is lifted since there are three different permittivities and the purely-uniaxial modes are not present. Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, in stark contrast to the situation in isotropic spheres. Fig. 3 (e) presents in the radiation patterns of representative biaxial-sphere modes.
We now analyze the high-order modes of uniaxial and biaxial ellipsoids in analogy to the anisotropic spheres. Similarly, we suggest that the anisotropic ellipsoid modes are specific modes or superpositions of isotropic ellipsoid modes and consider the modes expressed in cartesian coordinates (Niven’s functions). We apply our conditions with the second derivatives to these modes and observe that while the first type of Niven’s functions do not satisfy them, the second type do satisfy them and inside the inclusion correspond to the biaxial sphere modes [68, 69, 70, 71]. To impose continuity of for we project in the direction perpendicular to the interface with the unit vector and obtain for
Comparing inside the ellipsoid to the isotropic case we obtain that since and (inside and outside) are the same for this isotropic and anisotropic cases, we can equate to obtain the general (sum rule) resonance condition:
where [68]
For the sphere limit we get as expected. The modes in the form similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive (two conditions instead of one) and not likely to be realized experimentally. We proceed to the third-order mode of an anisotropic ellipsoid. We write the potential and field inside the inclusion Similarly, we equate inside the ellipsoid in the isotropic and anisotropic cases and obtain the resonance condition:
We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement In this way one can readily find the fourth-order biaxial ellipsoid mode and resonance condition.
Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes and the permittivity model used in Figs. 2,3 [54]. The sphere has a radius of and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of since we assume [67, 72, 31], and prove in the SM, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in . This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, as is explained in the Methods Section. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by where is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis. In this case, we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required. In Fig. 4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot for a dipole located at the same distance as the previous case and oriented along (31nm). As we predicted the modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles.
We then considered a biaxial subwavelength sphere composed of MoO3 with the permittivity model in Ref. [54]. The sphere is excited by a dipole located at with the dipole moment of We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL. Here, again, the dipole moment is directed along to preserve the angular dependency of the mode in the mode-excitation coefficient. In Fig. 5 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of . Crucially, our simulations confirm the resonance splitting effect predicted analytically, see arrows in Fig. 5. We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement. We also report very good agreement between the peak frequency widths obtained analytically and the simulations (accurate comparison would require a highly-dense frequency sweep and simulations that are very extensive).
Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with composed of MoO3 with the permittivity model in Ref. [54]. The ellipsoid is excited by a dipole located at with the dipole moment of We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL. In Fig. 6 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement. The predicted peaks were readily calculated by finding the maxima of . We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes (equivalent to Fig. 5) with excellent agreement. We note that the isotropic-ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following: As can be seen, for an isotropic ellipsoid with these semi axis ratios and and due to the anisotropic material these resonances are split, as can be seen in Fig. 6. In contrast to spheres, where the degeneracy is enforced by symmetry, here the degeneracy is accidental. Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellipsoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
Discussion
We derived the modes and eigenpermittivity relations of uniaxial and biaxial particles in the quasistatic regime. We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting. We also identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes. We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry. Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [73, 74, 75, 76, 32]. The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59]. Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [60, 77, 78]. Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
Acknowledgement
E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, and Y. Reches are acknowledged for the insightful comments. We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
I Methods
I.1 Correspondence between the derived modes and the radiation patterns
Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres. We utilize the TM vector spherical harmonics (VSH) modes for a general spherical particle given by where
and are the spherical bessel function and spherical Hankel function of the first kind and [79, 20]. Calculating the eigenpermittivity from these modes in the limit of gives the quasistatic eigenvalues. For example, the full electrodynamic mode gives the following eigenpermittivities for a sphere with nm and nm : which are very close to the quasistatic results of Even though these modes are usually considered for isotropic spheres, the results can be applied to anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs. We write a TM VSH outside the sphere:
Using the vector identity for we get for the first term:
For the second term we employ the same identity to get
Then, we utilize the spherical-harmonic identity [80]
to obtain
All in all we get:
We see that when since the overall scaling is of in agreement with the quasistatic analysis. For since and the first term, which is transverse and has scaling as expected, dominates. Thus, one can conclude from the superposition of that participate in the quasistatic anisotropic modes, that the radiation pattern will be of the form:
where the sum is over the mode orders which participate in the anisotropic-sphere modes. Please note that which gives overall scaling of the radiated field and we get the following radiation pattern:
One can readily see that the mode of a subwavelength sphere gives the dipole radiation pattern of . In addition, the radiation patterns of high-order anisotropic particle modes differ from the isotropic case. Note that while the uniaxial-sphere modes are identical to the isotropic sphere modes, in nearly all practical scenarios, their radiation patterns differ due to the fact that an excitation source will excite distinct superpositions of modes in the two cases.
I.2 Biaxial sphere modes as superpositions of spherical harmonics
We express the second-order biaxial modes as superpositions of spherical harmonics:
We now proceed to the third and fourth-order biaxial modes. We continue along the lines of the second-order biaxial mode and write Using the procedure above, it can be shown that and are the resonance conditions, respectively. Note that the modes and their resonance conditions readily follow.
I.3 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
I.3.1 Theory
We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the main text.
Dipole mode
The resonance condition is and we therefore find the maximum of which occurs for hBN [54] at The mode field inside the particle is given by
2nd order modes
For the first 2nd-order mode we get from the theory which for hBN approximately occurs at We add the degenerate modes, which are excited with equal strengths:
second mode
The derived resonance condition reads
For hBN this is approximately satisfied for Now we calculate the field by adding the degenerate modes:
3rd order modes
First mode
The resonance condition reads:
For hBN this occurs at . Adding the degenerate modes we get
To simplify things, we look at the plane
Second mode
The resonance condition is satisfied for hBN for Adding the modes we obtain:
We plot these field magnitude distributions in Fig. 4.
One can express the eigenvalue in the denominator in terms of the physical parameter using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs. 2 and 3). Note that the first sum includes only the dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper. It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.
Supplementary Material
Completeness of the solution
We now prove that the derived uniaxial-sphere modes are all the eigenmodes. Let us assume that there is an additional eigenstate besides the four modes for each . Since the isotropic-sphere modes span the space inside the sphere, this eigenstate is a superposition of the remaining isotropic-sphere modes
An eigenstate must satisfy anisotropic Laplace’s equation inside the inclusion, which reads:
Since it is a superposition of isotropic-sphere modes we write:
However, all these modes have orders of that are 2 or higher (including from ) and each function is fundamentally different from the others. Hence, such an eigenstate does not exist, and we have a contradiction. Conclusion: the derived uniaxial sphere modes are all the modes. Note that the uniaxial mode basis does not span space as in the standard case.
Similarly, we write for a biaxial particle
If we get which is a uniaxial medium, and we therefore exclude this case. All the modes are isotropic (from the boundary conditions) and the modes satisfy when Even if such modes exist, their axial permittivities must satisfy both Laplace’s equation (in a non-trivial manner) and the boundary condition, which imposes strong constraints. As a result, their potential contribution is typically weak. Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
Eigenmode expansion for an anisotropic inclusion
Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [79, 67]. Here we expand the quasielectrostatic potential for an anisotropic inclusion. For concreteness, we consider a uniaxial inclusion and the derived eigenstates. Generalizing it to a biaxial inclusion readily follows.
We start by writing Poisson’s equation for a medium with an anisotropic inclusion:
where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity. Now we move one term to the lhs in order to be able to define an eigenvalue equation.
where . The Green function in this case will be for a medium with an inclusion with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
where we defined to adjust it to an eigenvalue equation with an operator operating on and we performed integration by parts.
We express the electric potential and expand it using the anisotropic eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
Now we expand using the eigenstates of a particle with anisotropy in the axis:
where is the electric potential for the dipole in free space. All in all we get: